OpenStudy (anonymous):

change given equation into slope intercept form 5x+2y=10

OpenStudy (anonymous):

Hi.

OpenStudy (anonymous):

can u help me please give me answer or something

OpenStudy (anonymous):

If you write out the slope intercept formula, then i will.

OpenStudy (anonymous):

It looks something like y=m........

OpenStudy (anonymous):

Slope-Intercept form is y = mx + b Just solve for y. ~ 5x + 2y = 10 Subtract 5x. 2y = -5x + 10 Divide by 2. y = -5/2x + 5